OpenStudy (anonymous):

Find the center (h,k) and radius r of the circle 5x^(2)+5y^(2)+30x+20y-60=0

OpenStudy (campbell_st):

complete the square for x and y after dividing each term by 5 x^2 + 6x + y^2 + 4y = 12 (x^2 + 6x +9)+(y^2 +4y + 4) = 12 + 9 + 4 (x+3)^2 + (y + 2)^2 = 25 the circle has a centre at (-3, -2) and radius 5